# Deterministic Discrete Denoising

### v2

@@ 4.2 Word-Level Text Generation
Table 1: Evaluation of LM1B text samples by generative perplexity (Gen. PPL), entropy, and MAUVE score. The bottom two rows show our results using the UDLM model trained with the code from Schiff et al. (2025). The herding margin is set to $\delta=0.0005$ for $T=1024$ and $\delta=0.005$ for $T=128$. Best Gen. PPL values are indicated in bold. Published values from Schiff et al. (2025).
$T=1024$ $T=128$
Method Gen. PPL Entropy MAUVE Gen. PPL Entropy MAUVE
($\downarrow$) ($\uparrow$) ($\uparrow$) ($\downarrow$) ($\uparrow$) ($\uparrow$)
@@
UDLM (reproduced) 098.75 6.83 0.0116 098.33 6.81 0.0119
UDLM herding (ours) 046.71 6.30 0.0136 060.39 6.34 0.0135
4.3 Categorical-Valued Image Generation
We apply our herding-based algorithm to image generation, following the experiments of D3PM (Austin et al., 2021) and UDLM (Schiff et al., 2025). While pixel values are ordinal by nature, they are treated as discrete categorical variables for the purpose of evaluating generative performance. Namely, the forward uniform noising process assigns equal probability to transitions to neighboring and distant pixel values alike.
The dataset is discretized CIFAR-10 (Krizhevsky & Hinton, 2009), where the pixel intensities of the $32\times 32$ RGB images are discretized to 8-bit integer values. The model is based on a U-Net architecture (Ronneberger et al., 2015), which is identical to that used in Schiff et al. (2025). For evaluation, we computed Fréchet inception distance (FID) (Heusel et al., 2017) and inception score (IS) (Salimans et al., 2016) on 50,000 generated samples, using the PyTorch evaluation code  [^1] Available at https://github.com/w86763777/pytorch-image-generation-metrics with a pretrained Inception v3 model (Szegedy et al., 2016).
From the evaluation results shown in Table 2, we confirm that the trained unconditional UDLM model and conditional UDLM model (D-CFG with $\gamma=1$) reproduce the published values nearly identically. The proposed herding-based denoising consistently outperforms stochastic denoising in both FID and IS across all cases. Remarkably, the herding-based denoising with $T=128$ steps achieves performance comparable to that of stochastic denoising with $T=1024$ steps.
Table 2: Evaluation of CIFAR-10 image samples by FID and IS. The bottom two rows show our results using the UDLM models trained with the code from Schiff et al. (2025). D-CFG denotes conditional models with $\gamma=1$. The herding margin is set to $\delta=0.07$ for $T=128$ and $\delta=0.15$ for the other cases. Best values are indicated in bold. Published values from Austin et al. (2021) and Schiff et al. (2025).
$T=1000$ $T=1024$ (D-CFG) $T=128$ (D-CFG)
Method FID ($\downarrow$) IS ($\uparrow$) FID ($\downarrow$) IS ($\uparrow$) FID ($\downarrow$) IS ($\uparrow$)
@@ 5 Discussion
The deterministic denoising algorithm defines a mapping from the initial states $(\bm{x}^{(1:L)}_{T},\bm{w}^{(1:L)}_{T})$ to their final states $(\bm{x}^{(1:L)}_{0},\bm{w}^{(1:L)}_{0})$, which induces a flow of probability mass within the discrete state space $\mathcal{V}^{L}$. Each initial state encodes a generated sample $\bm{x}^{(1:L)}_{0}$, whose required precision is determined by the granularity of the piecewise isometric mapping. As our approach derandomizes Markov chains, the probability flow can be guided by the transition probabilities, as demonstrated in the numerical experiments with the conditional UDLM model (D-CFG) on the CIFAR-10 dataset. Our algorithm could also be combined with existing inference-time scaling methods (Kit et al., 2025; Singhal et al., 2025; Pani et al., 2025) by updating the particles deterministically.
The initial weight vectors $\bm{w}_{T}$ are currently chosen randomly from the uniform distribution on $0,1]^{K}$, which is a natural starting point as explained in Section A. Scaling the initial weights can influence the balance between the determinism and randomness of the algorithm. Moreover, there is no reason that all time steps should be treated equally in the accumulation of the discrepancy between the empirical and the target distributions. It is natural to consider weighted accumulation, where the weight scheduling can be an important factor in denoising performance.
@@
These observations on weighted accumulation and delayed switching naturally lead to the idea of continuous-time formulation. Discrete diffusion models formulated as continuous-time Markov chains (CTMCs) have been studied extensively (Campbell et al., 2022; Lou et al., 2024), and can also be naturally derandomized using the continuous-time variant of the herding system, which exhibits chaotic billiard dynamics (Suzuki, 2015). As an alternative, chaotic billiard sampling (Suzuki et al., 2013; Lee et al., 2025), which can be extended to multi-state spin models such as the Potts model (Suzuki, 2013), can also be applied. Compared with the CTMC formulation, these derandomized billiard systems require computing the exact state transition times induced by collisions with the billiard boundary, which is more time-consuming. Since the exact computation may not contribute to the performance, time-discretized computation is considered more practical. Nonetheless, continuous-time formulations may offer advantages in dedicated hardware implementations, as demonstrated in the integrated-circuit implementation of chaotic Boltzmann machines (Yamaguchi et al., 2019). This research direction is related to acceleration methods for the reverse processes of CTMCs (Park et al., 2025; Ren et al., 2025b).
As mentioned in the introduction, we focused on the uniform diffusion process in this paper. However, the proposed herding-based denoising algorithm can also be applied to other Markov-chain formulations of discrete diffusion models. While in Chen et al. (2024) the state transition time from the absorbing state is externally specified, the proposed deterministic algorithm may allow the transition time to be determined by the system itself according to the transition probability, which is conceptually related to an adaptive sampler for masked diffusion (Ben-Hamu et al., 2025). As the remasking mechanism, as employed in LLaDA (Nie et al., 2025; Zhu et al., 2025), makes the reverse process of masked diffusion more dynamic and enables iterative refinement, our approach could be applied to derandomize the remasking process to further enhance the refinement.
@@
Acknowledgments
This work was supported by ALCA-Next JPMJAN23F2 and Moonshot R&D JPMJMS2021 from the Japan Science and Technology Agency (JST), by Project JPNP14004 from the New Energy and Industrial Technology Development Organization (NEDO), and by KAKENHI 25K21300 from Japan Society for the Promotion of Science (JSPS).
Appendix A Bounded Dynamics of Herding-Based Denoising Algorithm
As a theoretical justification of our proposed algorithm in Section 3, we provide a proof of the boundedness of weight vector $\bm{w}_{t}$ and the $O(T^{-1})$ convergence of empirical sample distribution to the averaged probability vector.
$\sum_{i}w_{t-1,i}=\sum_{i}(w_{t,i}+p_{t-1,i}-x_{t-1,i})=\sum_{i}w_{t,i}+1-1=\sum_{i}w_{t,i}.$
Let $\mu=\sum_{i}w_{t,i}/K$ be the average weight. The initial weight satisfies $w_{T,i}\geq\mu-1$ because $w_{T,i}$ is chosen from $[0,1]$. Now assume that $w_{t,i}\geq\mu-1$ holds for some $t$. If $x_{t-1,i}=0$, then $w_{t-1,i}=w_{t,i}+p_{t-1,i}-x_{t-1,i}\geq w_{t,i}\geq\mu-1$. If $x_{t-1,i}=1$, then $w_{t,i}+p_{t-1,i}$ attains the maximum value, which is at least the average $\mu$. Hence, $w_{t-1,i}=w_{t,i}+p_{t-1,i}-x_{t-1,i}\geq\mu-1$. Thus, in both cases, we obtain $w_{t-1,i}\geq\mu-1$, and by induction the inequality holds for all $t$. This establishes a uniform lower bound for each component of $w_{t}$. Since a lower bound exists and the total weight $\sum_{i}w_{t,i}$ is conserved, every component of $\bm{w}_{t}$ must remain bounded from above as well, and therefore $\bm{w}_{t}$ is bounded. The same argument applies to the delayed-switching version, in which case the lower bound becomes $\mu-1-\delta$.
The $O(T^{-1})$ convergence follows from the boundedness of $\bm{w}_{t}$, as
$\frac{1}{T}\sum_{\tau=0}^{T-1}(\bm{p}_{\tau}-\bm{x}_{\tau})=\frac{1}{T}\sum_{\tau=0}^{T-1}(\bm{w}_{\tau}-\bm{w}_{\tau+1})=\frac{1}{T}(\bm{w}_{0}-\bm{w}_{T})\to 0\qquad(T\to\infty).$
For initialization, it is natural to choose $\mathbf{w}_{T}$ randomly from the uniform distribution on $0,1]^{K}$ for the following reason. The stationary distribution of the uniform-diffusion Markov chain is the uniform distribution over the sample space $\mathcal{V}$. Therefore, we regard $\bm{w}_{T}$ as a weight vector obtained by iterating the herding updates for the uniform probability vector $\bm{p}_{t}=\bm{1}/K$ for all $t\geq T$, as if the updates had continued from $t=\infty$. Then, after a transient period, each weight $w_{t,i}$ increases by $1/K$ at each iteration but decreases by $1$ once every $K$ iterations. As a result, $w_{T,i}$ stays within an interval of length $1$. Hence, we choose the uniform distribution on $0,1]^{K}$ as a natural starting point.

### v3

@@ 3.2 Bounded Dynamics, Convergence, and Initialization
Table 2: Evaluation of LM1B text samples by generative PPL (Gen. PPL), entropy, and MAUVE score. The bottom two rows show our results using the UDLM model trained with the code from (Schiff et al., 2025). The herding margin is set to $\delta=0.0005$ for $T=1024$ and $\delta=0.005$ for $T=128$. Best Gen. PPL values are indicated in bold. Published values from (Schiff et al., 2025).
Method $T=1024$ $T=128$
Gen. PPL ($\downarrow$) Entropy ($\uparrow$) MAUVE ($\uparrow$) Gen. PPL ($\downarrow$) Entropy ($\uparrow$) MAUVE ($\uparrow$)
@@
UDLM (reproduced) 098.75 6.83 0.424 098.33 6.81 0.369
UDLM herding (ours) 046.71 6.30 0.648 060.39 6.34 0.473
Table 3: Evaluation of CIFAR-10 image samples by FID and IS. The bottom two rows show our results using the UDLM models trained with the code from (Schiff et al., 2025). D-CFG denotes conditional models with $\gamma=1$. The herding margin is set to $\delta=0.07$ for $T=128$ and $\delta=0.15$ for the other cases. Best values are indicated in bold. Published values from (Austin et al., 2021) and (Schiff et al., 2025).
Method $T=1000$ $T=1024$ (D-CFG) $T=128$ (D-CFG)
FID ($\downarrow$) IS ($\uparrow$) FID ($\downarrow$) IS ($\uparrow$) FID ($\downarrow$) IS ($\uparrow$)
@@
Table 4: Evaluation of TSP solutions obtained by DIFUSCO. The tour lengths and performance gaps, averaged over all instances, are shown for the TSP-500, TSP-1000, and TSP-10000 datasets. The performance gap is computed relative to the best-known baselines (Fu et al., 2021). Greedy and Sampling denote the generation strategies based on one sample with $T=50$ steps and 16 samples with $T=10$ steps, respectively. 2-opt indicates post-processing using the 2-opt local search heuristic. The herding margin is set to $\delta=0.2$. Best values for each strategy are indicated in bold. Published values of DIFUSCO (Sun & Yang, 2023).
Strategy Refinement TSP-500 TSP-1000 TSP-10000
Length ($\downarrow$) Gap ($\downarrow$) Length ($\downarrow$) Gap ($\downarrow$) Length ($\downarrow$) Gap ($\downarrow$)
Best-known 16.55 00.00% 23.12 00.00% 71.78 00.00%
Greedy 18.35 26.14 98.15
Greedy (reproduced) 18.24 10.24% 25.61 10.80% 97.49 35.82%
Greedy + herding (ours) 18.06 09.14% 25.94 12.21% 93.18 29.82%
Greedy 2-opt 16.80 23.56 73.99
Greedy (reproduced) 2-opt 16.81 01.63% 23.56 01.91% 73.94 03.01%
Greedy + herding (ours) 2-opt 16.73 01.12% 23.58 02.01% 73.31 02.13%
Sampling 17.23 25.19 95.52
Sampling (reproduced) 17.16 03.69% 24.69 06.80% 94.65 31.86%
Sampling + herding (ours) 17.07 03.19% 24.77 07.16% 92.29 28.58%
Sampling 2-opt 16.65 23.45 73.89
Sampling (reproduced) 2-opt 16.65 00.66% 23.38 01.13% 73.99 03.08%
Sampling + herding (ours) 2-opt 16.63 00.52% 23.37 01.11% 73.56 02.49%
4 Experiments
We demonstrate the effectiveness of the proposed discrete denoising algorithm through experiments on text and image generation and diffusion-based combinatorial optimization. For clarity of presentation, we reproduce the trained models of UDLM (Schiff et al., 2025) and DIFUSCO (Sun & Yang, 2023), and evaluate the improvements that result solely from introducing the herding algorithm into the reverse denoising process. All experiments are conducted using the official codes provided by the authors. Our algorithm is implemented with only about 20 lines of code modifications, which are included in the Appendix for reproducibility.
4.1 Character-Level Text Generation
We first evaluate the proposed algorithm in character-level text generation, using the text8 dataset (Mahoney, 2011). The vocabulary size is $35$ and the output length is fixed to $256$ characters. The model architecture follows the UDLM paper (Schiff et al., 2025), which is based on a Transformer with 92.4M parameters. For evaluation, we generate 1,024 samples for each condition and compute the generative perplexity (PPL) based on the GPT-2 Large pretrained model (Radford et al., 2019) and the entropy.
Figure 1 shows the generative PPL and entropy of the generated samples. The herding-based denoising with $\delta=0$ consistently outperforms stochastic denoising in terms of PPL, achieving up to a $10\times$ improvement for $T=1024$ inference steps, albeit at the expense of entropy. Increasing the margin $\delta$ interpolates between these two extremes, allowing us to improve PPL with a moderate trade-off in entropy.
As DNDM (Chen et al., 2024) provides a deterministic denoising algorithm for masked diffusion, it is informative to contrast its performance on the text8 dataset with ours, although the experimental settings are not exactly identical. DNDM reduces the PPL from 1465.75 of a 1000-step D3PM absorb to 600.02 using 256-step inference. In contrast, our method reduces the PPL from 459.53 of a 1024-step UDLM to 79.16 using 256-step herding, indicating stronger performance in both the magnitude of improvement and the achieved generative PPL.
4.2 Word-Level Text Generation
We also evaluate the proposed algorithm in word-level text generation, using the one billion words dataset (LM1B) (Chelba et al., 2013). The vocabulary size is $30,522$ and the output length is $128$ tokens. The model architecture follows the UDLM paper (Schiff et al., 2025), which is based on a Transformer with 139M parameters. For evaluation, we generate 1,024 samples for each condition and compute the generative PPL and MAUVE score (Pillutla et al., 2021) based on the GPT-2 Large pretrained model (Radford et al., 2019) as well as the entropy.
Table 2 shows the evaluation results. In our reproduction experiment, the UDLM model trained using wrapped LM1B sequences performs worse than the published results (Schiff et al., 2025). Nevertheless, the herding-based denoising substantially reduces the PPL compared to stochastic denoising for both $T=1024$ and $T=128$ steps, outperforming the published UDLM (uniform) and MDLM (masked) results and even the autoregressive (AR) model. We also observe slightly lower entropy values and improved MAUVE scores. Overall, the proposed algorithm significantly improves the coherence and linguistic quality of the generated text while affecting the distributional metrics only minimally. It should also be noted that, unlike stochastic denoising, the generative PPL improves as the number of inference steps $T$ increases, which suggests that iterative refinement is enhanced through sample generation following the reverse Markov chain more faithfully.
4.3 Categorical-Valued Image Generation
We apply our herding-based algorithm to image generation, following the experiments of D3PM (Austin et al., 2021) and UDLM (Schiff et al., 2025). While pixel values are ordinal by nature, they are treated as discrete categorical variables in this experiment. Specifically, the forward uniform noising process assigns equal probability to transitions to neighboring and distant pixel values alike, which results in suboptimal performance compared to other noising processes. Accordingly, this experiment is intended only to demonstrate the improvements due to the derandomization.
We use the discretized CIFAR-10 dataset (Krizhevsky & Hinton, 2009), where the pixel intensities of the $32\times 32$ RGB images are discretized to 8-bit integer values. The model is based on a U-Net architecture (Ronneberger et al., 2015), which is identical to that used in (Schiff et al., 2025). For evaluation, we computed Fréchet inception distance (FID) (Heusel et al., 2017) and inception score (IS) (Salimans et al., 2016) on 50,000 generated samples, using the PyTorch evaluation code  [^1] Available at https://github.com/w86763777/pytorch-image-generation-metrics with a pretrained Inception v3 model (Szegedy et al., 2016).
From the evaluation results in Table 3, we confirm that the trained unconditional UDLM model and conditional UDLM model (D-CFG with $\gamma=1$) reproduce the published values nearly identically. The herding-based denoising consistently outperforms stochastic denoising in both FID and IS across all cases. Remarkably, the herding-based denoising with $T=128$ steps achieves performance comparable to that of stochastic denoising with $T=1024$ steps.
4.4 Diffusion-Based Combinatorial Optimization
As an alternative evaluation approach, we apply the herding-based denoising to combinatorial optimization. DIFUSCO (Sun & Yang, 2023) is a framework for solving combinatorial optimization problems using discrete diffusion models, where discrete diffusion with Bernoulli noise has been shown to perform significantly better than continuous diffusion with Gaussian noise. In the Bernoulli case, the generative process can be derandomized using the herding-based algorithm, and its performance can be quantitatively evaluated using the objective values of the generated solutions.
We conduct experiments on traveling salesman problems (TSPs), following the experimental settings in Sun & Yang (2023). Solutions are obtained from the discrete diffusion model with binary variables ($K=2$), where the prediction model is based on anisotropic graph neural networks (AGNNs). We use the official code and pretrained checkpoints provided by the authors (Sun & Yang, 2023) to evaluate the improvement by the herding-based denoising.
Table 4 shows the average objective values of the generated TSP solutions. These results show that our experiments reproduce the published values very closely. The proposed herding-based denoising outperforms stochastic denoising in almost all cases.
@@ 5 Discussion
The proposed denoising algorithm defines a mapping from the initial states $(\bm{x}^{(1:L)}_{T},\bm{w}^{(1:L)}_{T})$ to their final states $(\bm{x}^{(1:L)}_{0},\bm{w}^{(1:L)}_{0})$, which induces a flow of probability mass within the discrete state space $\mathcal{V}^{L}$. Each initial state encodes a generated sample $\bm{x}^{(1:L)}_{0}$, whose required precision is determined by the granularity of the piecewise isometric mapping. As our approach derandomizes Markov chains, the probability flow can be guided by the transition probabilities, as demonstrated in the numerical experiments with the conditional UDLM model (D-CFG) on the CIFAR-10 dataset. Our algorithm could also be combined with existing inference-time scaling methods (Kit et al., 2025; Singhal et al., 2025; Pani et al., 2025) by updating the particles deterministically.
The initial weight vectors $\bm{w}_{T}$ are currently drawn randomly from the uniform distribution on $0,1]^{K}$, which is a natural starting point as explained in Section 3.2. Scaling the initial weights can influence the balance between the determinism and randomness of the algorithm. Moreover, there is no reason that all time steps should be treated equally in the accumulation of the discrepancy between the empirical and the target distributions. It is therefore natural to consider weighted accumulation, which has close connections to the Frank–Wolfe algorithm (Bach et al., 2012) and Bayesian quadrature (Huszár & Duvenaud, 2012). The weight scheduling can be an important factor in denoising performance.
@@
These observations on weighted accumulation and delayed switching naturally lead to the idea of continuous-time formulation. Discrete diffusion models formulated as continuous-time Markov chains (CTMCs) have been studied extensively (Campbell et al., 2022; Lou et al., 2024), and can also be naturally derandomized using the continuous-time variant of the herding system, which exhibits chaotic billiard dynamics (Suzuki, 2015). As an alternative, chaotic billiard sampling (Suzuki et al., 2013; Lee et al., 2025), which can be extended to multi-state spin models such as the Potts model (Suzuki, 2013), can also be applied. Compared with the CTMC formulation, these derandomized billiard systems require computing the exact state transition times induced by collisions with the billiard boundary, which is more time-consuming. Since the exact computation may not contribute to the performance, a time-discretized version, closely related to our proposed algorithm, is considered more practical. Nonetheless, continuous-time formulations may offer advantages in dedicated hardware implementations, as demonstrated in the integrated-circuit implementation of chaotic Boltzmann machines (Yamaguchi et al., 2019). This research direction is related to acceleration methods for the reverse processes of CTMCs (Park et al., 2025; Ren et al., 2025b).
As mentioned in the introduction, we focused on the uniform diffusion process in this paper. However, the proposed herding-based denoising algorithm can also be applied to other Markov-chain formulations of discrete diffusion models. While in (Chen et al., 2024) the state transition times from the absorbing state are externally specified, the proposed deterministic algorithm may allow the transition times to be determined by the system itself according to the transition probability, which is conceptually related to an adaptive sampler for masked diffusion (Ben-Hamu et al., 2025). As the remasking mechanism, as employed in LLaDA (Nie et al., 2025; Zhu et al., 2025), makes the reverse process of masked diffusion more dynamic and enables iterative refinement, our approach could be applied to derandomize the remasking process to further enhance the refinement.
@@
Impact Statement
This paper proposes a novel derandomization approach for discrete diffusion models. While generative models may involve broad societal impacts depending on their use, we do not foresee any specific negative societal impacts arising directly from our contribution.
Appendix A Experimental Details
A.1 Code Modifications for UDLM
We present below the code diff that integrates the proposed herding-based denoising algorithm into the official UDLM codebase (Schiff et al., 2025), for reproducibility of our experimental results in Sections 4.1, 4.2 and 4.3.
--- a/diffusion.py
+++ b/diffusion.py
@@ -142,6 +142,8 @@ class Diffusion(L.LightningModule):
self.lr = self.config.optim.lr
self.sampling_eps = config.training.sampling_eps
+ self.deterministic_sampling = getattr(self.config.sampling, ’name’, ’gumbel’) == ’deterministic’
+ self.deterministic_switch_threshold = getattr(self.config.sampling, ’deterministic_switch_threshold’, 1.0)
self.softplus = torch.nn.Softplus()
self.neg_infinity = -1_000_000.0
@@ -170,6 +172,22 @@ class Diffusion(L.LightningModule):
self._validate_configuration()
+ def _sample_categorical_stateful(self, categorical_probs, xt=None):
+ if not self.deterministic_sampling:
+ return _sample_categorical(categorical_probs)
+ self.errors += categorical_probs.to(self.errors.dtype)
+ xs = self._delayed_argmax(xt)
+ self.errors -= F.one_hot(xs, num_classes=self.vocab_size).to(self.errors.dtype)
+ return xs
+ def _delayed_argmax(self, xt):
+ if xt is None:
+ return self.errors.argmax(dim=-1)
+ errors_xmax, xmax = self.errors.max(dim=-1)
+ errors_xt = torch.gather(self.errors, -1, xt.unsqueeze(-1)).squeeze(-1)
+ delay_flag = (errors_xmax - errors_xt) < self.deterministic_switch_threshold
+ return torch.where(delay_flag, xt, xmax)
def _validate_configuration(self):
assert not (self.change_of_variables
and self.importance_sampling)
@@ -1123,6 +1141,9 @@ class Diffusion(L.LightningModule):
self.config.model.length
).to(self.device)
+ if self.deterministic_sampling:
+ self.errors = torch.rand(*xt.shape, self.vocab_size, device=xt.device, dtype=torch.float32)
timesteps = torch.linspace(
1, eps, self.config.sampling.steps + 1, device=self.device)
dt = (1 - eps) / self.config.sampling.steps
@@ -1243,7 +1264,7 @@ class Diffusion(L.LightningModule):
f"Diffusion type {self.diffusion} not implemented.")
# Sample from posterior
- xs = _sample_categorical(q_xs)
+ xs = self._sample_categorical_stateful(q_xs, xt)
if self.diffusion == ’absorbing_state’:
copy_flag = (xt != self.mask_index).to(torch.bool)
q_xs[copy_flag] = 0.0
@@ -1331,7 +1352,7 @@ class Diffusion(L.LightningModule):
f"Diffusion type {self.diffusion} not implemented.")
# Sample from posterior
- xs = _sample_categorical(q_xs)
+ xs = self._sample_categorical_stateful(q_xs, xt)
if self.diffusion == ’absorbing_state’:
copy_flag = (xt != self.mask_index).to(torch.bool)
q_xs[copy_flag] = 0.0
@@ -1449,7 +1470,7 @@ class Diffusion(L.LightningModule):
guided_probs = guided_log_probs.softmax(dim=-1)
# Sample from guided posterior
- xs = _sample_categorical(guided_probs)
+ xs = self._sample_categorical_stateful(guided_probs, xt)
if self.diffusion == ’absorbing_state’:
xs = torch.where(copy_flag.to(bool), xt, xs)
return xs, guided_probs, {’log_x_theta’: log_x_theta,
@@ -1561,7 +1582,7 @@ class Diffusion(L.LightningModule):
raise NotImplementedError(
f"Diffusion type {self.diffusion} not implemented.")
- xs = _sample_categorical(guided_probs)
+ xs = self._sample_categorical_stateful(guided_probs, xt)
if self.diffusion == ’absorbing_state’:
xs = torch.where(copy_flag, xt, xs)
A.2 Code Modifications for DIFUSCO
We present below the code diff that replaces the original denoising algorithm with the proposed herding-based algorithm in the official DIFUSCO codebase (Sun & Yang, 2023), for reproducibility of our experimental results in Section 4.4.
--- a/pl_tsp_model.py
+++ b/pl_tsp_model.py
@@ -134,8 +134,19 @@
else:
x0_pred_prob = x0_pred.reshape((1, points.shape[0], -1, 2)).softmax(dim=-1)
- xt = self.categorical_posterior(target_t, t, x0_pred_prob, xt)
- return xt
+ prob = self.categorical_posterior(target_t, t, x0_pred_prob, xt)
+ self.wt += prob.to(self.wt.dtype)
+ xs = self._delayed_argmax(xt)
+ self.wt -= F.one_hot(xs, num_classes=2).to(self.wt.dtype)
+ return xs
+ def _delayed_argmax(self, xt):
+ if xt is None:
+ return self.wt.argmax(dim=-1)
+ wp_max, x_max = self.wt.max(dim=-1)
+ wp_stay = torch.gather(self.wt, -1, xt.unsqueeze(-1)).squeeze(-1)
+ delay_flag = (wp_max - wp_stay) < self.args.ddd_margin # margin from run-time option
+ return torch.where(delay_flag, xt, x_max)
def gaussian_denoise_step(self, points, xt, t, device, edge_index=None, target_t=None):
with torch.no_grad():
@@ -202,6 +213,8 @@
steps = self.args.inference_diffusion_steps
time_schedule = InferenceSchedule(inference_schedule=self.args.inference_schedule,
T=self.diffusion.T, inference_T=steps)
+ self.wt = torch.rand(*xt.shape, 2).to(device) # initialize
# Diffusion iterations
for i in range(steps):
--- a/pl_meta_model.py
+++ b/pl_meta_model.py
@@ -136,14 +136,8 @@
sum_x_t_target_prob += x_t_source_prob_new[..., 1] * x0_pred_prob[..., 1]
- if target_t > 0:
- xt = torch.bernoulli(sum_x_t_target_prob.clamp(0, 1))
- else:
- xt = sum_x_t_target_prob.clamp(min=0)
- if self.sparse:
- xt = xt.reshape(-1)
- return xt
+ prob = sum_x_t_target_prob.clamp(0, 1).reshape(-1)
+ return torch.stack([1-prob, prob], dim=-1)
def gaussian_posterior(self, target_t, t, pred, xt):
"""Sample (or deterministically denoise) from the Gaussian posterior for a given time step.